# DiffusionNAG: Predictor-guided Neural Architecture Generation with Diffusion Models

### 31 Dec 2023 v2

@@ 3.3 The Effectiveness of Score Network for Neural Architectures
3.4 Adaptation Across Different Objectives
Table 5: Adaptation ability across different objectives.
Stats. Guidance
$f_{\phi_{\text{Clean}}}$ $f_{\phi_{\text{APGD}}}$
$f_{\phi_{\text{Blur}}}$
Test Accuracy Clean
Max
93.90
80.40
86.60
Mean
93.17
69.24
74.23
APGD
Max
2.60
26.60
22.50
Mean
1.51
17.66
14.02
Blur
Max
37.10
53.50
56.90
Mean
33.28
40.46
47.88
@@ 4 Related Work
Diffusion models, as demonstrated in prior work (Song & Ermon, 2019; Ho et al., 2020; Song et al., 2021b), are designed to reverse the data perturbation process, enabling them to generate samples from noisy data. They have achieved success in a variety of domains, including images (Nichol et al., 2022; Rombach et al., 2022), audio (Jeong et al., 2021; Kong et al., 2021), and graphs (Niu et al., 2020b; Jo et al., 2022). However, existing diffusion models are not well-suited for Neural Architecture Generation (NAG) because their primary focus is on unconditionally generating undirected graphs. To overcome this limitation, this study introduces a conditional diffusion-based generative framework tailored for generating architectures represented as directed acyclic graphs that meet specified conditions, such as accuracy requirements.
@@ 5 Conclusion
This study introduced a novel approach called DiffusionNAG, which is the paradigm shift from existing NAS methods. With the guidance of a predictor for a given task, DiffusionNAG can efficiently generate task-optimal architectures. Additionally, the introduction of a score network ensures the generation of valid architectures. Extensive experiments demonstrated that DiffusionNAG outperforms existing NAS, especially effective in the extensive search space. For reproducibility, we include “NAS Best Practices Checklist” in  Appendix A and the code in the supplementary file.
@@ Appendix A NAS Best Practice Checklist
      Did you release the code for the training pipeline used to evaluate the final architectures? We released the code for the training pipeline along with the supplementary file.
@@
      Did you release the code for the search space? In our main experiments, we utilized the NAS-Bench-201 search space, which is publicly available along with its description and code.
@@
      Did you release the hyperparameters used for the final evaluation pipeline, as well as random seeds? We use the NAS-Bench-201 pipeline and hyperparameters as our final evaluation pipeline. It can be found in the released code.
@@
      Did you release code for your NAS method? The code for our NAS method can be found in the supplementary file.
@@ B.2 Diffusion Process
Our approach aims to apply continuous-time generative diffusion processes to neural architectures. Specifically, we utilize the Variance Exploding (VE) SDE (Song et al., 2021b) to model the forward diffusion process for our neural architecture generation framework. The following SDE describes the process of the VE SDE (Jo et al., 2022):
@@ C.7 Implemetation Details
In Equation 11, which describes the diffusion process in DiffusionNAG, we set the minimum diffusion variance, denoted as $\sigma_{\min}$, to 0.1 and the maximum diffusion variance, denoted as $\sigma_{\max}$, to 5.0. These values determine the range of diffusion variances used during the diffusion process. Additionally, we utilize 1000 diffusion steps. This means that during the diffusion process, we iterate 1000 times to progressively generate and update the architectures. Furthermore, we set the sampling epsilon $\epsilon$ to $1\times 10^{-5}$. This value is used in the sampling step to ensure numerical stability (Song et al., 2021b).
@@ D.2 Improving Existing Bayesian Optimization NAS across Various Acquisition Functions
D.3 Analysis of Generated Neural Architectures across Different Objectives
@@
In the Adaptation Across Different Objectives part of  Section 3.4 in the main paper, we demonstrate the capability of our conditional generative framework to easily adapt to new tasks. This adaptability is achieved by seamlessly swapping the predictor, trained with the specific target objective, without the need to retrain the score network.
@@
D.4 Generated Neural Architectures Details
In this section, we provide detailed information on the top-5 architectures selected by the dataset-aware predictor following the generation process. The specific architectural details of the generated neural architectures can be found in Figure 8.
(a) CIFAR-10
(b) CIFAR-100
(c) Aircraft
(d) Oxford-IIIT Pets
Figure 8: Architecture Details in the NAS-Bench-201 Search Space. We generate 500 architectures and select 30 architectures using the dataset-aware surrogate model. Among these selected architectures, we provide the details of the architectures that achieve a top-5 accuracy for each dataset. Furthermore, we present the average accuracy for each dataset below the architecture structure, which is computed from 3 different seeds.
@@ D.5 Adaptation Across Different Objectives: Corruptions
In Table 5 of the main paper, we evaluate the adaptation abilities of DiffusionNAG to new tasks by swapping the task-specific predictor in a plug-and-play manner without retraining the score network. In the Section, we conduct further experiments to explore a broader range of corruptions introduced not only in glass blur but also in various corruptions in the NB201-based robustness benchmark (Jung et al., 2023). We conduct experiments to find robust neural architectures for 15 diverse corruptions at severity level 2 in CIFAR-10-C. Similar to Table 5 of the main paper, we demonstrated the capabilities of DiffusionNAG to easily adapt to new tasks without retraining the score network. We achieved this by replacing task-specific predictors in a plug-and-play manner. Specifically, we train a clean predictor, $f_{\bm{\phi}_{\text{clean}}}$), to predict the clean accuracy of neural architectures using randomly sampled pairs of architectures and their corresponding CIFAR-10 clean accuracy from the NB201-based robustness benchmark. Afterward, we train the target corruption predictor, $f_{\bm{\phi}_{\text{corruption}}}$, to predict the robust accuracy of neural architectures under specific corruption types. We create the training dataset by including pairs of architectures and their respective robust accuracy for each corruption scenario from the NB201-based robustness benchmark. With the guidance of the trained $f_{\bm{\phi}_{\text{clean}}}$, we generated a pool of 20 architectures ($S_{\text{clean}}$. Similarly, with the guidance of the task-specific predictors ($f_{\bm{\phi}_{\text{corruption}}}$), we generate another pool of 20 architectures for each target corruption ($S_{\text{corruption}}$). Subsequently, we retrieve the robust accuracy on both $f_{\bm{\phi}_{\text{clean}}}$-guided and $f_{\bm{\phi}_{\text{corruption}}}$-guided architecture pools ($S_{\text{clean}}$ and $S_{\text{corruption}}$) from the NB201-based robustness benchmark for each corruption scenario.
@@
D.6 Search strategy for the guidance strength
Table 11: Validity Results - Coarse Grid Search
Guidance Strength $10^{6}$ $10^{5}$ $10^{4}$
Validity 0.0 0.0 100.0
To effectively define the search range for guidance strength, we employ a following strategy:
Coarse grid search: Inspired by Lee et al. (2023), we first perform a coarse grid search on the guidance strength within a log scale, identifying a strength where validity becomes stable. For instance, within the NB201 search space under the Transferable NAS scenario (Table 11), a guidance strength of $10^{4}$ yielded 100.0% valid neural architectures.
Fine-grained grid search: Building upon the coarse grid search, we perform a more fine-grained grid search around the point identified in the coarse grid search phase. This involves incrementally adjusting the guidance strength until we identify the optimal point where validity remains intact.
Subsequently, we employ a balanced exploitation-exploration sampling strategy that iteratively decreases the guidance strength from this identified point through the fine-grained grid search. Influenced by the analysis of guidance strength and fidelity/diversity in previous studies (Dhariwal & Nichol, 2021; Ho & Salimans, 2022; Marwood et al., 2023), we designed this strategy to balance the trade-off between exploitation and exploration: Higher guidance strength focuses more on the modes of the guiding predictor, prioritizing fidelity over diversity, which is similar to exploitation. On the contrary, utilizing lower guidance strength reduces dependence on the guiding predictor, enabling increased engagement in exploration.
Table 12: Ablation Study - Guidance Strength Distribution of generated neural architectures across various sampling strategies. All methods generated neural architectures within the same sampling time.
Stats. Fixed at a large value Balanced exploitation-exploration (Ours) Fixed at a small value
Avg. Acc. of Top-5 (%) 93.262 94.156 93.978
Standard Deviation (%) 0.395 0.658 0.953
Ablation studies on the impact of guidance strength.
To empirically verify how this guidance strength affects the performance and further provide insights for future follow-up studies, we compared the neural architecture distributions generated using our sampling strategy (balanced exploitation-exploration sampling strategy), both when the guidance scale was fixed at a large value and when it was fixed at a small value (Please see Table 12). We conducted experiments on the CIFAR-10 dataset within the NB201 search space under the Transferable NAS scenario. The fixed large and small values correspond to the maximum and minimum guidance strengths in our strategy of gradually decreasing guidance strength with a scale 5000, respectively (i.e., 15000, 10000, 5000). As we mentioned in Section 3.1 in the main paper, after generating neural architectures by each sampling strategy, we rank the generated architecture pools using a meta-learned dataset-aware predictor and also provide the average accuracy of the Top-5 neural architectures.
As evident from Table 12, the average accuracy of Top-5 neural architectures generated with the Fixed at a large value strategy (93.262%) is lower than the average accuracy of Top-5 neural architectures (94.156%) achieved by Ours, possibly due to heavily engaging in exploitation. On the other hand, fixing the guidance strength to one small value allows for greater engagement in exploration, leading to the generation of neural architectures with more diversity. However, since it guides towards regions far from the target sub-region, the average accuracy of the generated Top-5 neural architectures (93.978%) is lower than the result obtained by Ours, i.e., when the guidance strength is 0, it can be considered equivalent to random sampling. Given the trade-off between exploration and exploitation depending on the guidance strength, we progressively adjust it during the generation of neural architectures. This sampling approach offers the advantage of balancing exploitation-exploration by gradually reducing the guidance strength.
Pitfalls of solely relying on a fixed large guidance strength.
Using only one fixed large guidance strength to generate neural architectures may not lead to the generation of an optimal neural architecture. In other words, utilizing only one large guidance strength can lead to a heavy reliance on the guiding predictor (exploitation), which may result in phenomena corresponding to model collapse. However, by adopting our aforementioned sampling strategy that gradually decreases the guidance strength for generating neural architectures, we can achieve a balance between exploitation and exploration. As the guidance strength diminishes, the bias towards the predictor decreases, mitigating the risk of model collapse. Empirically, as seen in Table R2, the Standard Deviation values for the accuracy of the set of neural architectures generated by the fixed at a large value method and the proposed sampling method are 0.395 and 0.658, respectively. The fixed at a large value method tends to produce relatively less diverse neural architecture structures. As a result, the proposed method (Top-5 accuracy: 94.154%) outperforms the fixed at a large value method (Top-5 accuracy: 93.262%).
Appendix E Future Work
In this section, we explore potential future research directions using DiffusionNAG. Firstly, we consider employing a discrete diffusion model as our primary framework, departing from the continuous diffusion model utilized in the main paper. Then, we discuss the feasibility of generating optimal generative models using the DiffusionNAG framework based on the discrete diffusion model. Finally, we discuss the application of DiffusionNAG as a post-training structured pruning approach to design sparse foundation models.
E.1 DiffusionNAG to design optimal generative models
We explore the feasibility of generating optimal generative models using a discrete diffusion model under the proposed DiffusionNAG framework. While there are a lot of design choices to generate diffusion models, we consider transformer-based generative models. Inspired by Wang et al. (2020), a NAS work focusing on designing vanilla transformers (although not generation model-based NAS work), we can define a search space that encompasses key architectural design choices, such as the number of transformer blocks, the number of MLP layers within each block, the number of attention layers within each block, and the dimension of hidden states in attention or MLP layers. Similar to a set of visual tokens (Chang et al., 2022) or vocab tokens in the NLP field (Devlin et al., 2019), we can create a set of architectural tokens by representing all combinations of architectural design choices as architectural tokens (i.e., one-hot vectors). Similar to dividing a vacant canvas into image patches and predicting visual tokens for each patch (Chang et al., 2022), in NAS, we can consider that the macro skeleton of architectures (the entire graph) corresponds to the vacant canvas, where nodes of the graph correspond to image patches. Subsequently, we perform architectural token predictions for each node in the graph. To train the score model, our approach involves sampling graphs from the search space, partially masking nodes, and predicting architectural tokens for the masked nodes. This transforms the problem into a cross-entropy minimization task, predicting probabilities for each masked token and computing the cross-entropy between the ground truth one-hot token and the predicted token. The performance metric in this scenario may vary depending on the target generative task. For instance, we can consider FID or IS as performance metrics for image synthesis. We can train the predictor guiding the architecture generation process of DiffusionNAG to predict the performance of an architecture for the target generative task. By utilizing the trained predictor, we can guide the generation of DiffusionNAG to generate an optimal architecture of a generative model for the target task.
E.2 DiffusionNAG to design sparse foundation models
To explore efficient sparse architectures for foundation models, we can leverage DiffusionNAG, similar to the research idea discussed in Section E.1. Applying diffusion model to generate diffusion models. Among various potential directions, one illustrative example is to consider DiffusionNAG as a post-training structured pruning approach designed to induce sparsity in Large Language Models (LLMs). This post-training structured pruning method starts from pre-trained models and eliminates redundant architecture elements along with their parameters. We find this approach reasonable, particularly when considering the impracticality of iteratively pre-training or iteratively pruning models at the scale of LLMs with hundreds of billions of parameters. Specifically, since the majority of LLMs are transformer-based models, we can explore sparse architectures within the search space that was defined in our response to Section E.1. Our goal would be to identify the optimal architecture that minimizes the accuracy drops within a resource budget (e.g., memory footprint or inference time). In other words, we aim to find the best architecture that enhances inference speed or reduces memory footprint without compromising accuracy. We can address this problem by employing the proposed DiffusionNAG method. First, we can train the score model of DiffusionNAG in a manner similar to that explained in (Tan et al., 2019). Secondly, we can formulate a multi-objective function to evaluate the performance of sparse architecture candidates. For example, we may consider existing multi-objectives proposed by ProxylessNAS (Cai et al., 2019) or MNASNet (Tan et al., 2019) to find an optimal balance between accuracy and latency. Using the multi-objective function, we can train the predictor to predict the performance of given architecture candidates. Guided by the trained predictor, DiffusionNAG can generate an optimal sparse architecture (graph). In the final step, we selectively activate only the architectural components and parameters of LLMs that align with the generated graph. In other words, we prune (deactivate) architectural components such as blocks, MLP layers, and attention layers that do not correspond to the graph. This process results in a resource-efficient model, enhancing the sparsity of the model and contributing to improved efficiency in terms of inference latency or memory utilization.

### 24 Mar 2024 v4

@@ 4 Related Work
Diffusion models, as demonstrated in prior work (Song & Ermon, 2019; Ho et al., 2020; Song et al., 2021), are designed to reverse the data perturbation process, enabling them to generate samples from noisy data. They have achieved success in a variety of domains, including images (Nichol et al., 2022; Rombach et al., 2022), audio (Jeong et al., 2021; Kong et al., 2021), and graphs (Niu et al., 2020b; Jo et al., 2022). However, existing diffusion models are not well-suited for Neural Architecture Generation (NAG) because their primary focus is on unconditionally generating undirected graphs. To overcome this limitation, this study introduces a conditional diffusion-based generative framework tailored for generating architectures represented as directed acyclic graphs that meet specified conditions, such as accuracy requirements.
@@ 5 Conclusion
This study introduced a novel conditional Neural Architecture Generation (NAG) framework called DiffusionNAG, which is the paradigm shift from existing NAS methods by leveraging diffusion models. With the guidance of a property predictor for a given task, DiffusionNAG can efficiently generate task-optimal architectures. Additionally, the introduction of a score network ensures the generation of valid neural architectures. Extensive experiments under two key predictor-based NAS scenarios demonstrated that DiffusionNAG outperforms existing NAS methods, especially effective in the large search space. We believe that our success underscores the potential for further advancements in NAS methodologies, promising accelerated progress in the development of optimal neural architectures.
Reproducibility Statement
To ensure reliable and reproducible results, we have provided the source code, available at https://github.com/CownowAn/DiffusionNAG, and the “NAS Best Practices Checklist” in Appendix A. Furthermore, we have described the details of the diffusion process and models used in DiffusionNAG in Appendix B, as well as the experimental details in Appendix C.
Ethics Statement
This paper introduces a novel conditional Neural Architecture Generation framework called DiffusionNAG, aimed at advancing the field of generative models, especially for neural architectures. We believe that our work can open up new avenues for neural architecture search/generation, achieving significant acceleration and improved search performance in extensive experimental settings. Although the possibility of causing any negative societal impacts through our work is low, we should be cautious about the potential for our work to be utilized to generate neural architectures for unethical purposes.
Acknowledgement
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2023-00256259), Center for Applied Research in Artificial Intelligence (CARAI) grant funded by DAPA and ADD (UD190031RD), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2022-0-00713), and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2019-0-00075, Artificial Intelligence Graduate School Program(KAIST)).
@@ Appendix A NAS Best Practice Checklist
      Did you release the code for the training pipeline used to evaluate the final architectures? We released the code for the training pipeline.
@@
      Did you release the code for the search space? In our main experiments, we utilized the NAS-Bench-201 and MobileNetV3 search space, which is publicly available along with its description and code.
@@
      Did you release the hyperparameters used for the final evaluation pipeline, as well as random seeds? Yes, it can be found in the code we provide.
@@
      Did you release code for your NAS method? The code for our work is available on GitHub: https://github.com/CownowAn/DiffusionNAG.
@@ B.2 Diffusion Process
Our approach aims to apply continuous-time generative diffusion processes to neural architectures. Specifically, we utilize the Variance Exploding (VE) SDE (Song et al., 2021) to model the forward diffusion process for our neural architecture generation framework. The following SDE describes the process of the VE SDE (Jo et al., 2022):
@@ C.7 Implemetation Details
In Equation 11, which describes the diffusion process in DiffusionNAG, we set the minimum diffusion variance, denoted as $\sigma_{\min}$, to 0.1 and the maximum diffusion variance, denoted as $\sigma_{\max}$, to 5.0. These values determine the range of diffusion variances used during the diffusion process. Additionally, we utilize 1000 diffusion steps. This means that during the diffusion process, we iterate 1000 times to progressively generate and update the architectures. Furthermore, we set the sampling epsilon $\epsilon$ to $1\times 10^{-5}$. This value is used in the sampling step to ensure numerical stability (Song et al., 2021).
@@ D.2 Improving Existing Bayesian Optimization NAS across Various Acquisition Functions
D.3 Adaptation Across Different Objectives
Table 9: Adaptation ability across different objectives.
Stats. Guidance
$f_{\phi_{\text{Clean}}}$ $f_{\phi_{\text{APGD}}}$
$f_{\phi_{\text{Blur}}}$
Test Accuracy Clean
Max
93.90
80.40
86.60
Mean
93.17
69.24
74.23
APGD
Max
2.60
26.60
22.50
Mean
1.51
17.66
14.02
Blur
Max
37.10
53.50
56.90
Mean
33.28
40.46
47.88
D.4 Analysis of Generated Neural Architectures across Different Objectives
@@
In the Adaptation Across Different Objectives part of  Section D.3, we demonstrate the capability of our conditional generative framework to easily adapt to new tasks. This adaptability is achieved by seamlessly swapping the predictor, trained with the specific target objective, without the need to retrain the score network.
@@ D.5 Adaptation Across Different Objectives: Corruptions
In Table 9, we evaluate the adaptation abilities of DiffusionNAG to new tasks by swapping the task-specific predictor in a plug-and-play manner without retraining the score network. In the Section, we conduct further experiments to explore a broader range of corruptions introduced not only in glass blur but also in various corruptions in the NB201-based robustness benchmark (Jung et al., 2023). We conduct experiments to find robust neural architectures for 15 diverse corruptions at severity level 2 in CIFAR-10-C. Similar to Table 9, we demonstrated the capabilities of DiffusionNAG to easily adapt to new tasks without retraining the score network. We achieved this by replacing task-specific predictors in a plug-and-play manner. Specifically, we train a clean predictor, $f_{\bm{\phi}_{\text{clean}}}$), to predict the clean accuracy of neural architectures using randomly sampled pairs of architectures and their corresponding CIFAR-10 clean accuracy from the NB201-based robustness benchmark. Afterward, we train the target corruption predictor, $f_{\bm{\phi}_{\text{corruption}}}$, to predict the robust accuracy of neural architectures under specific corruption types. We create the training dataset by including pairs of architectures and their respective robust accuracy for each corruption scenario from the NB201-based robustness benchmark. With the guidance of the trained $f_{\bm{\phi}_{\text{clean}}}$, we generated a pool of 20 architectures ($S_{\text{clean}}$. Similarly, with the guidance of the task-specific predictors ($f_{\bm{\phi}_{\text{corruption}}}$), we generate another pool of 20 architectures for each target corruption ($S_{\text{corruption}}$). Subsequently, we retrieve the robust accuracy on both $f_{\bm{\phi}_{\text{clean}}}$-guided and $f_{\bm{\phi}_{\text{corruption}}}$-guided architecture pools ($S_{\text{clean}}$ and $S_{\text{corruption}}$) from the NB201-based robustness benchmark for each corruption scenario.